OpenStudy (anonymous):

Use a graph of the function to explain the geometric meaning of the value of the integral.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

\[\int\limits_{0}^{3} (x ^{3}-6x) dx\]

OpenStudy (anonymous):

thats the integral

OpenStudy (michele_laino):

we have to draw the graph of the function x^3-6x

OpenStudy (michele_laino):

here is my screenshot:

OpenStudy (michele_laino):

the x-intercepts of that function, are: x^3-6x=0 what is x?

OpenStudy (michele_laino):

\[\Large \begin{gathered} {x^3} - 6x = 0 \hfill \\ x\left( {{x^2} - 6} \right) = 0 \hfill \\ \end{gathered} \]